OpenStudy (anonymous):

3x-y=9, find the inverse

OpenStudy (anonymous):

3y-x=9, solve for y, 3y=x+9, y=(x+9)/3, or y=1/3x+3.

geerky42 (geerky42):

To find the inverse, switch x and y to each other then isolate y.

OpenStudy (anonymous):

\[f(x)=3x-9 \rightarrow f^{-1}(x)=3y-9 \implies x=3y-9 \rightarrow y^{-1}=\frac{ x }{ 3 }+3\]

OpenStudy (anonymous):

@ahekker

OpenStudy (anonymous):

Thank you, ahekker.